For what integral values of ‘a’ , the equation x2-x(1-a) – (a+2) = 0 has integral roots. Find the roots.

Grade:11

1 Answers

Ajay
7 years ago
For roots to be a interger the D should be perfect square.
Here D = (1-a)+4(a+2) = a2+2a+9
Now D can take up any up the follwing values to be a perfect square
1,4,9,16.25 etc
selecting D =9 will give you inegral values for a as well as roots.
so  a2+2a+9 = 9, solve for a and get the remaining answer